ThisGirlPretty:

What is the equivalent of pi over 3 radians in degrees? 30 degrees 45 degrees 60 degrees 90 degrees

ThisGirlPretty:

@Vocaloid

Rainluke01:

60

Shadow:

@Rainluke01 Please explain how you got your answer.

Rainluke01:

I went to google and typed pi/3 radians in degrees

Rainluke01:

and it sent me here https://www.quia.com/jg/321176list.html

Shadow:

Well, we want TGP to be able to do it on her own, in the event of a test for example :) We have, \[\frac{ \pi }{ 3 }\] One radian is equal to \[\frac{ 180 }{ \pi }\] Therefore we do \[\frac{ \pi }{ 3 } \times \frac{ 180 }{ \pi } = 60 degrees\]

Shadow:

The reason you use the value for radians is so you can cancel out pi (or radians) to be left with degrees. The conversion is real simple as all you have to do is flip pi over 180 to meet each case respectively. For example, if we wanted to convert 60 degrees back to radians One degree is equal to \[\frac{ \pi }{ 180 }\] So we do, \[60 \times \frac{ \pi }{ 180 } = \frac{ \pi }{ 3 }\]

ThisGirlPretty:

Yes thank you